OpenStudy (anonymous):

Which of the following words does not indicate multiplication? A. times B. of C. product D. more than

OpenStudy (anonymous):

d?

OpenStudy (ryan123345):

more than = addition

OpenStudy (anonymous):

D. More than

OpenStudy (anonymous):

Three times the sum of a number and seven is two less than five times the number. A. 3x + 7 = 5x – 2 B. 3(x + 7) = 5x – 2 C. 3(x + 7) = 2 – 5x D. 3x + 7 = 2 – 5x

OpenStudy (anonymous):

c

OpenStudy (anonymous):

The quotient of a number and four is twelve less than the number. What is the number? A. –4 B. 9.6 C. 8 D. 16